OpenStudy (anonymous):

Can somebody please help me on one problem? What is 4x+2y=20 in slope intercept form?

OpenStudy (imstuck):

What do you have?

OpenStudy (anonymous):

ok,what's ur problem?

OpenStudy (anonymous):

4x+2y=20 in slope intercept form?

OpenStudy (anonymous):

help anyone

OpenStudy (anonymous):

well just keep 2y in a side and then subtract 4x from both sides,it would be : 2y = -4x - 20 and the standard form is y = mx + b in which m = slope and b = y-intercept. and now you have to divide both sides by 2,then it would be: y = -2x - 10

OpenStudy (anonymous):

now it's slope intercept form ;)

OpenStudy (anonymous):

u sure its not positive 10?

OpenStudy (anonymous):

oh yes sry ;) i made a mistake yes it's y = -2x + 10

OpenStudy (anonymous):

any problem?

OpenStudy (anonymous):

lol its okay

OpenStudy (anonymous):

fine ;)

OpenStudy (anonymous):

can u help on something else? @PFEH.1999

OpenStudy (anonymous):

ofcourse ;)

OpenStudy (anonymous):

The kite has vertices D(0, 3b), E(a, 0), and F(0, -5b). What are the coordinates of G?

OpenStudy (anonymous):

@PFEH.1999

OpenStudy (anonymous):

sry

OpenStudy (anonymous):

?

OpenStudy (anonymous):

well,i think you should find the length of DG,DE,EF,FG

OpenStudy (anonymous):

i did lol

OpenStudy (anonymous):

\( DG + EF = DE + GF \)

OpenStudy (anonymous):

(-a,0)?

OpenStudy (anonymous):

you should spare me a moment...i'll check it out.

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

correct ;)

OpenStudy (anonymous):

THANKKSSSS

OpenStudy (anonymous):

ur welcome ;)